OpenStudy (anonymous):

6.15(x – 8.86) = 21.83 A. 12.8 B. 12.7 C. 12.6 D. 11.7 Which is the best estimate of x based on rounding the constants and coefficients in the equation to the nearest integer?

OpenStudy (anonymous):

6(x - 9) = 22 6x - 54 = 22 6x = 22 + 54 6x = 76 x = 76/6 = 12.666 rounds to 12.7